Two forces act on a point objects as follows :100N at 170° and 100N at 50°.find their resultant
1
Expert's answer
2019-10-08T10:05:41-0400

The resultant force of two forces is a vector sum of the vectors of these forces. A vector sum can be found by the law of cosines:

"F_{resultant}^2=F_1^2+F_2^2+2F_1F_2cos(\alpha-\beta)"


After setting given values we have:

"F_{resultant}=\sqrt{10^4+10^4+2*10^2*10^2*(-1/2)}"


"F_{resultant}=\sqrt{10^4+10^4-10^4}"

"F_{resultant}=\sqrt{10^4}"

"F_{resultant}=100N"


The action angle of the resultant force equals:

"\gamma=(\alpha+\beta)/2=(170°+50°)/2=110°"


Answer: The resultant force equals 100N at 110°.